# Impact of cavities on the detection of quadratically coupled ultra-light dark matter

### v1

@@
Impact of cavities on the detection of quadratically coupled ultra-light dark matter
Clare Burrage clare.burrage@nottingham.ac.uk    Angus Macdonald angus.macdonald@nottingham.ac.uk School of Physics and Astronomy, University of Nottingham, University Park, Nottingham NG7 2RD, UK    Michael P. Ross mpross2@uw.edu    Gray Rybka grybka@uw.edu University of Washington, Seattle, WA 98195, USA    Elisa Todarello elisa.todarello@nottingham.ac.uk School of Physics and Astronomy, University of Nottingham, University Park, Nottingham NG7 2RD, UK
?abstractname?
@@ I Introduction
Galactic and cosmological scale observations, particularly in the face of the robustness of General Relativity [1], have motivated the existence of an unobserved “dark matter” constituting around $25\%$ of the energy balance of the Universe [2]. Light scalars have been considered as an origin of this phenomenon [3, 4, 5]: a class of feebly interacting particle, with mass $\ll 1\mbox{ eV}$ Such scalars are common objects in models of physics beyond the Standard Model, appearing in theories beyond General Relativity [6, 7, 8], and models attempting to resolve fine tuning in the Standard Model such as the QCD axion [9, 10].
Light scalars must be produced non-thermally in the early universe to be oscillating in their potential today, and thus behave as cold dark matter. Example of such production mechanisms are vacuum realignment [11, 12, 13, 14], decay of heavy fields such as moduli [15] or decay of topological defects [16]. If the field mass is $\ll 1$ eV, typical occupation numbers are so high that it is possible to model the dark matter in our Galaxy as an incoherent superposition of classical scalar waves [17] [^1] See Refs [18, 19, 20, 21] for a discussion of quantum corrections to the classical field behaviour.. In order to detect or constrain such scalar dark matter fields, it is necessary to understand how these particles interact with matter. Absent a principle to forbid such interactions, we expect them to be present, and they may be universal, for example: arising from a conformal rescaling of the metric, they may be mediated through a Higgs portal coupling, or arise directly from interactions between the scalar and fundamental particles. Key work by Damour and Donoghue, [22, 23], showed that if such a light scalar evolves with time, it induces time dependent shifts in effective constants of physics - particularly in the fine structure constant, fermionic masses and the QCD confinement scale, and how the couplings to fundamental particles can be expressed as couplings to atoms. This framework has been used to predict scalar-mediated phenomena such as ‘$5^{\rm th}$ forces’, Equivalence Principle violating effects, and shifts to atomic spectra. These have in turn been used to justify searches for such particles through means such as tests of the universality of free fall, and frequency shifts in atomic clocks [24, 25, 26, 27, 28, 29, 30, 31, 32]. Such theories will also be constrained with future experimental advances, including with atom interferometry [33, 34, 35, 36, 37], hyperfine transitions [38, 39], molecular spectroscopy [40] and nuclear clocks [41].
@@ II Quadratically coupled scalar dark matter
The constant $\alpha$, has dimensions $[\mbox{mass}]^{-2}$, and parametrises the interaction strength between scalar and matter. When matter is non-relativistic a coupling of the scalar field to energy density can arise from a universal coupling to matter, such as a conformal coupling [6, 7, 8], or through a coupling of the scalar to the Higgs boson  [55, 56, 45, 57, 58, 59, 60, 61], which when the Higgs is integrated out leaves an effective coupling of the form in equation (1) at low energies. It can also arise from coupling the scalar directly to the fundamental bosons and gluons [62, 63, 47, 42, 54, 43], in which case composition dependence may arise, discussed further in section II.1. In this work we are interested in the leading order effects of a universal coupling to matter. If the coupling to matter is composition dependent, the effects discussed in this work might vary somewhat in strength for different materials, but will otherwise remain. We assume that the $\rho\alpha\phi^{2}$ is the leading order interaction between the scalar field and matter, but allow for the possibility that higher order interactions may be present. To ensure that such higher order terms remain subdominant we will require $|\alpha\phi^{2}|\ll 1$.
@@
From equation (2), it is apparent that a local energy density can shift the effective (squared) mass of the scalar field, $m_{eff}^{2}=m^{2}+\rho\alpha$. We note that when $\rho\alpha<-m^{2}$ the effective squared mass may be negative. For the largest values of $\alpha$ that we will consider, everyday materials can lead to large increases in the mass of a scalar that is otherwise extremely light in vacuum, leading to Compton wavelengths on the order of the size of experimental apparatus:
@@ II.1 Connection to ‘dilaton coefficients’
$\mathcal{L}_{int}=\frac{(\kappa\phi)^{2}}{2}\left[\frac{d_{e}}{4}F_{\mu\nu}F^{\mu\nu}-\frac{d_{g}\beta_{3}}{2g_{3}}F^{A}_{\mu\nu}F^{A\leavevmode\nobreak\ {\mu\nu}}-\sum_{i={u,d,e}}(d_{m_{i}}+\gamma_{m_{i}}d_{g})m_{i}\bar{\psi}_{i}{\psi}_{i}\right]\;,$ (4)
with $\kappa=\sqrt{4\pi G}$. $F$ and $F^{A}$ are photon and gluon field strength tensors respectively, $g_{3}$ is the QCD gauge coupling, which runs according to $\beta_{3}$, $\gamma_{m_{i}}$ provides the anomalous dimension responsible for the running of quark masses, and $d_{e},\leavevmode\nobreak\ d_{g},\leavevmode\nobreak\ d_{m_{i}}$ are the dimensionless ‘dilaton coefficients’. These interactions induce an effective scalar dependence in the fundamental constants of nature - namely the fine structure constant, fermionic masses, and the QCD mass scale:
$\begin{split}\alpha_{EM}(\phi)&=\alpha_{EM}\left(1+d_{e}\frac{(\kappa\phi)^{2}}{2}\right)\\ m_{i}(\phi)&=m_{i}\left(1+d_{m_{i}}\frac{(\kappa\phi)^{2}}{2}\right)\\ \Lambda_{3}(\phi)&=\Lambda_{3}\left(1+d_{g}\frac{(\kappa\phi)^{2}}{2}\right)\end{split}$ (5)
@@
$m_{A}(\phi)=m_{A}\left(1+\tilde{\alpha}_{A}\frac{(\kappa\phi)^{2}}{2}\right)$ (6)
@@
$\begin{split}d_{\hat{m}}&=\frac{m_{u}d_{m_{u}}+m_{d}d_{m_{d}}}{m_{u}+m_{d}}\\ d_{\delta m}&=\frac{m_{u}d_{m_{u}}-m_{d}d_{m_{d}}}{m_{u}-m_{d}}\end{split}$ (8)
@@ II.2 Homogeneous Background
$\rho_{DM}=\frac{1}{2}m^{2}\phi_{\infty}^{2}\leavevmode\nobreak\ .$ (12)
Throughout this work, we will take the local dark matter energy density today to be $\rho_{DM}\simeq 0.4\leavevmode\nobreak\ \mbox{GeV}\leavevmode\nobreak\ \mbox{cm}^{-3}$ [64].
@@ III.1 Spherically Symmetric Uniform Object
for constant $\rho_{S}$. We assume this sphere is isolated, such that at an infinite distance from the object, the field tends to the limit of the of the homogeneous background. This allows us to factorise the field into homogeneous and spatiality varying components:
@@
$\nabla^{2}\varphi_{S}(r)=\alpha\rho(r)\varphi_{S}(r)\leavevmode\nobreak\ .$ (16)
@@
$\varphi_{S}(r)=\begin{cases}A^{in}_{S}\frac{\cosh(kr)}{r}\ +B^{in}_{S}\frac{\sinh(kr)}{r}&r<R\\ A^{out}_{S}+\dfrac{B^{out}_{S}}{r}&r>R\leavevmode\nobreak\ .\\ \end{cases}$ (17)
@@
$\varphi_{S}(r)=\begin{cases}\operatorname{sech}(kR)\frac{\sinh(kr)}{kr}&r<R\\ 1+\dfrac{q(kR)R}{r}&r>R\end{cases}$ (18)
@@
$q(kR)=\frac{\tanh(kR)-kR}{kR}\leavevmode\nobreak\ .$ (19)
@@
$q(kR)\approx\begin{cases}-\frac{1}{3}(kR)^{2}&|kR|\ll 1\\ -1&|kR|\gg 1\end{cases}$ (20)
The case $|kR|\ll 1$ is interpreted as a regime of weak coupling between sphere and field, in which the characteristic length scale of the field in matter, $1/|k|$, is considerably larger than the sphere. Hence, in this region, the field is minimally perturbed, and $\varphi_{S}\approx 1$. Conversely, in the case $|kR|\gg 1$, the sphere and field can be considered strongly coupled. As can be seen in Fig. (1), this leads to a general trend in which the field is increasingly screened in the vicinity of the sphere as $|kR|$ increases. However, in the case $\alpha<0$, we observe additional phenomenology. Since $k\in i\mathbb{R}$, Eq. (19) contains a term $\sim\tan(|k|R)$, the field diverges asymptotically at $|k|R=\pi\left(n+\frac{1}{2}\right)$, for integer $n$, and is hence enhanced in the neighbourhood of these points. Whilst the condition $|\alpha\phi^{2}|\ll 1$ does not hold sufficiently close to these divergences, there may still be $|k|$ for which some degree of enhancement can be physical, depending on the value of $m$ associated with the field. Conversely, we note that there are additionally values for which $\tan(|k|R)=|k|R$, and so the effective scalar charge of the source vanishes, $q(kR)=0$, and the field profile around the sphere will be indistinguishable from the uncoupled case.
@@
$\varphi_{\oplus}(r)=\varphi_{S}(r\;;R_{\oplus},\leavevmode\nobreak\ k_{\oplus})\;,$ (21)
for which $k_{\oplus}=\sqrt{\rho_{\oplus}\alpha}$, $\rho_{\oplus}\simeq 5510\leavevmode\nobreak\ \mbox{kg m}^{-3}$ and $R_{\oplus}=6.4\times 10^{6}\leavevmode\nobreak\ m$.
?figurename? 1: Evaluation of the field profile at the surface of a solid sphere radius $R$, varying $|\alpha\rho_{S}R^{2}|=|kR|^{2}$. The top panel illustrates the case $\alpha>0$, in which the field at the surface is screened for $\alpha=1/\rho R^{2}$. The lower panel demonstrates the case $\alpha<0$. Maxima and minima indicate the locations of field divergences and zeros respectively. We note that the magnitude of the field is displayed given that, in the $\alpha<0$ case, sign changes occur between adjacent extrema.
@@ III.2 Spherical Cavity
where, $r,\leavevmode\nobreak\ \theta$ form part of the spherical co-ordinate basis about the cavity centre, defining $\theta\in\{0\leq\theta<\pi\}$ such that $\theta=0$ aligns with the direction of the field gradient. A caveat here is that, under our own description, the background field to the cavity should tend to the homogeneous solution away from the matter sourcing it. This requirement is in conflict with Eq. (24), which diverges as $r\rightarrow\infty$. However, we may discard this problem if the background field is sufficiently slowly varying, such that it can be described only by the leading order terms in its Taylor expansion in the vicinity of the cavity. These terms can be used to define $a_{0},\leavevmode\nobreak\ a_{1}$. For example, by expanding $\varphi_{\oplus}(r^{\prime};k_{\oplus},R_{\oplus})$, about the position of the cavity relative to the centre of the Earth, $\vec{r}_{c}$, we might define $a_{0},\leavevmode\nobreak\ a_{1}$ using the field sourced by the Earth:
@@
The validity of this interpretation when evaluating $a_{0},\leavevmode\nobreak\ a_{1}$ for different parameters is discussed in Sec. IV. With the inclusion of these boundary conditions, the system is no longer spherically symmetric and the equation of motion for the field is:
$\frac{1}{r^{2}}\partial_{r}(r^{2}\partial_{r}\varphi_{c})+\frac{1}{r^{2}\sin\theta}\partial_{\theta}\left(\sin\theta\partial_{\theta}\varphi_{c}\right)=\rho\alpha\varphi_{c}$ (26)
By imposing the boundary conditions in Eq. (24), demanding continuity and differentiability of the field at the boundaries $r\in\{0,\leavevmode\nobreak\ R_{1},\leavevmode\nobreak\ R_{2}\}$, and through orthogonality of angular modes, solutions simplify to the form:
$\varphi_{c}(r)=\begin{cases}A_{0}^{in}+A_{1}^{in}r\cos\theta&r<R_{1}\\ \\ \begin{split}&\left(A_{0}^{walls}[j_{0}(-ikr)]+B_{0}^{walls}[(i)y_{0}(-ikr)]\right)\\ &+\left(A_{1}^{walls}[(-i)j_{1}(-ikr)]+B_{1}^{walls}[(-1)y_{1}(-ikr)]\right)\cos\theta\end{split}&R_{1}<r<R_{2}\\ \\ \left(a_{0}+B_{0}^{out}r^{-1}\right)+\left(a_{1}r+B_{1}^{out}r^{-2}\right)\cos\theta&r>R_{2}\end{cases}$ (27)
Where $A^{in/out}_{l},\leavevmode\nobreak\ B^{in/out}_{l}$ are constants of integration, $j_{l}(x),\leavevmode\nobreak\ y_{l}(x)$ are spherical Bessel functions of the first and second kind respectively, $k=\sqrt{\alpha\rho_{c}}$ and $i$ is the imaginary unit. Further detail on the origin of these terms, and analytical expressions for these coefficients are are presented in Appendix A.
@@
$a_{0}\rightarrow A^{in}_{0}$ $=a_{0}\times\frac{1}{kR_{1}\sinh(k(R_{2}-R_{1}))+\cosh(k(R_{2}-R_{1}))}$ (28)
$a_{1}\rightarrow A^{in}_{1}$ $=a_{1}\times\frac{3kR_{2}}{3\sinh(k(R_{2}-R_{1}))+3kR_{1}\cosh(k(R_{2}-R_{1}))+(kR_{1})^{2}\sinh(k(R_{2}-R_{1}))}$ (29)
@@
?figurename? 2: The modification of coefficients describing the field inside a spherical cavity: $R_{1}=10\leavevmode\nobreak\ \mbox{cm}$, $R_{2}=11\leavevmode\nobreak\ \mbox{cm}$ and $\rho_{c}=2700\leavevmode\nobreak\ \mbox{kg m}^{-3}$. Panels in the top row relate to coefficients for the positive $\alpha$ case, and the bottom row the negative $\alpha$ case. The left panels give the magnitude of the ratio between the constant component of the field profile inside the cavity, as in (27), and the boundary condition in Eq. (24). The right gives the same but for the linear component of the profile and boundary conditions. Note that the magnitude is plotted since for the $\alpha<0$ case, components of the field profile inside the cavity may change sign with respect to the boundary conditions. The blue dashed line indicates where $\alpha=k_{c}^{2}/\rho_{c}$.
@@
We demonstrate the effect of this modification in Fig. 2 for a small test cavity, inner radius $R_{1}=10\leavevmode\nobreak\ \mbox{cm}$ and outer radius $R_{2}=11\leavevmode\nobreak\ \mbox{cm}$, and with density similar to aluminium $\rho_{c}=2700\leavevmode\nobreak\ \mbox{kg m}^{-3}$. These dimensions have been chosen as to approximate the scale and properties of laboratory scale apparatus such as optical cavities. Other choices of size and material will exhibit qualitatively similar behaviour.
@@
As with the uniform spherical case, we may choose to broadly divide the solution for the field into two regimes, by comparing the value of $|k|$ to the thickness of the wall of the cavity. The critical value for a cavity is: $k_{c}=1/\left(R_{2}-R_{1}\right)$. For $|\alpha|\ll k_{c}^{2}/\rho_{c}$, such that the wavelength of the field in the cavity walls is larger than the cavity thickness $R_{2}-R_{1}$, Fig. 2 shows that the cavity has minimal impact on the amplitude of the field in the interior. Indeed, Eqs. (28), (29) demonstrate that $A_{0},\leavevmode\nobreak\ A_{1}$ should be well approximated by $a_{0},\leavevmode\nobreak\ a_{1}$ for $|k|\ll k_{c}$, aside for the region $1/R_{1}\lesssim|k|<k_{c}$, in which polynomial decay arising from the $kR_{1}$ terms in the denominators could be observed.

### v2

@@
Abstract
@@ I Introduction
Galactic and cosmological scale observations, particularly in the face of the robustness of General Relativity [77], have motivated the existence of an unobserved “dark matter” constituting around $25\%$ of the energy balance of the Universe [4]. Light scalars have been considered as an origin of this phenomenon [46, 55, 6]: a class of feebly interacting particle, with mass $\ll 1\mbox{ eV}$. Such scalars are common objects in models of physics beyond the Standard Model, appearing in theories beyond General Relativity [25, 30, 56], and models attempting to resolve fine tuning in the Standard Model such as the QCD axion [75, 76].
Light scalars must be produced non-thermally in the early universe to be oscillating in their potential today, and thus behave as cold dark matter. Examples of such production mechanisms are vacuum realignment [1, 37, 65, 7], decay of heavy fields such as moduli [24] or decay of topological defects [33]. If the field mass is $\ll 1$ eV, typical occupation numbers are so high that it is possible to model the dark matter in our Galaxy as an incoherent superposition of classical scalar waves [45] [^1] See Refs [68, 28, 38, 39] for a discussion of quantum corrections to the classical field behaviour.. In order to detect or constrain such scalar dark matter fields, it is necessary to understand how these particles interact with matter. Absent a principle to forbid such interactions, we expect them to be present, and they may be universal, for example: arising from a conformal rescaling of the metric, they may be mediated through a Higgs portal coupling, or arise directly from interactions between the scalar and fundamental particles. Key work by Damour and Donoghue, [31, 32], showed that if such a light scalar evolves with time, it induces time dependent shifts in effective constants of physics - particularly in the fine structure constant, fermionic masses and the QCD confinement scale, and how the couplings to fundamental particles can be expressed as couplings to atoms. This framework has been used to predict scalar-mediated phenomena such as $5^{\rm th}$ forces’, Equivalence Principle violating effects, and shifts to atomic spectra. These have in turn been used to justify searches for such particles through means such as tests of the universality of free fall, and frequency shifts in atomic clocks [59, 63, 53, 44, 8, 69, 41, 73, 20]. Such theories will also be constrained with future experimental advances, including with atom interferometry [2, 21, 40, 3, 9], hyperfine transitions [52, 57], molecular spectroscopy [62] and nuclear clocks [15].
@@ II Quadratically coupled scalar dark matter
The constant $\alpha$, has dimensions $\mbox{mass}]^{-2}$, and parametrises the interaction strength between scalar and matter. When matter is non-relativistic a coupling of the scalar field to energy density can arise from a universal coupling to matter, such as a conformal coupling [25, 30, 56], or through a coupling of the scalar to the Higgs boson [22, 66, 64, 5, 16, 43, 42, 23], which when the Higgs is integrated out leaves an effective coupling of the form in equation (1) at low energies. It can also arise from coupling the scalar directly to the fundamental bosons and gluons [54, 58, 17, 19, 18, 49, 29], in which case composition dependence may arise, discussed further in section II.1. In this work we are interested in the leading order effects of a universal coupling to matter. If the coupling to matter is composition dependent, the effects discussed in this work might vary somewhat in strength for different materials, but will otherwise remain. We assume that the $\rho\alpha\phi^{2}$ is the leading order interaction between the scalar field and matter, but allow for the possibility that higher order interactions may be present. To ensure that such higher order terms remain subdominant we will require $|\alpha\phi^{2}|\ll 1$.
@@
From equation (2), it is apparent that a local energy density can shift the effective (squared) mass of the scalar field, $m_{eff}^{2}=m^{2}+\rho\alpha$. We note that when $\rho\alpha<-m^{2}$ the effective squared mass of a field oscillating about $\phi=0$ may be negative. As will be discussed in later sections, this may cause the expectation value of the field expectation to roll to a new minimum, which can have important implications with regards to the stability of field solutions, and the validity of the scalar theory in Eq. (1) [13, 54, 12, 10, 11, 48]. For the largest values of $\alpha$ that we will consider, everyday materials can lead to large increases in the mass of a scalar that is otherwise extremely light in vacuum, leading to Compton wavelengths on the order of the size of experimental apparatus:
@@ II.1 Connection to ‘dilaton coefficients’
$\mathcal{L}_{int}=\frac{(\kappa\phi)^{2}}{2}\left[\frac{d_{e}}{4}F_{\mu\nu}F^{\mu\nu}-\frac{d_{g}\beta_{3}}{2g_{3}}F^{A}_{\mu\nu}F^{A\penalty 10000\ {\mu\nu}}-\sum_{i={u,d,e}}(d_{m_{i}}+\gamma_{m_{i}}d_{g})m_{i}\bar{\psi}_{i}{\psi}_{i}\right]\;,$ (4)
with $\kappa=\sqrt{4\pi G}$. $F$ and $F^{A}$ are photon and gluon field strength tensors respectively, $g_{3}$ is the QCD gauge coupling, which runs according to $\beta_{3}$, $\gamma_{m_{i}}$ provides the anomalous dimension responsible for the running of quark masses, and $d_{e},\penalty 10000\ d_{g},\penalty 10000\ d_{m_{i}}$ are the dimensionless ‘dilaton coefficients’. These interactions induce an effective scalar dependence in the fundamental constants of nature - namely the fine structure constant, fermionic masses, and the QCD mass scale:
$\begin{split}\alpha_{EM}(\phi)&=\alpha_{EM}\left(1+d_{e}\frac{(\kappa\phi)^{2}}{2}\right)\;,\\ m_{i}(\phi)&=m_{i}\left(1+d_{m_{i}}\frac{(\kappa\phi)^{2}}{2}\right)\;,\\ \Lambda_{3}(\phi)&=\Lambda_{3}\left(1+d_{g}\frac{(\kappa\phi)^{2}}{2}\right)\;.\end{split}$ (5)
@@
$m_{A}(\phi)=m_{A}\left(1+\tilde{\alpha}_{A}\frac{(\kappa\phi)^{2}}{2}\right)\;,$ (6)
@@
$\begin{split}d_{\hat{m}}&=\frac{m_{u}d_{m_{u}}+m_{d}d_{m_{d}}}{m_{u}+m_{d}}\;,\\ d_{\delta m}&=\frac{m_{u}d_{m_{u}}-m_{d}d_{m_{d}}}{m_{u}-m_{d}}\;.\end{split}$ (8)
@@ II.2 Homogeneous Background
$\rho_{DM}=\frac{1}{2}m^{2}\phi_{\infty}^{2}\penalty 10000\ .$ (12)
Throughout this work, we will take the local dark matter energy density today to be $\rho_{DM}\simeq 0.4\penalty 10000\ \mbox{GeV}\penalty 10000\ \mbox{cm}^{-3}$ [34].
@@ III.1 Spherically Symmetric Uniform Object
for constant $\rho_{S}$. We assume this sphere is isolated, such that at an infinite distance from the object, the field tends to the limit of the of the homogeneous background. This allows us to factorise the field into homogeneous and spatially varying components:
@@
$\nabla^{2}\varphi_{S}(r)=\alpha\rho(r)\varphi_{S}(r)\penalty 10000\ .$ (16)
@@
Here, $k$ acts as growth rate/wavenumber for the field inside the sphere, given as $k=\sqrt{\rho_{S}\alpha}$. Note that when $\alpha$ is positive: $k\in\mathbb{R}$ and the field solutions are hyperbolic, but when $\alpha$ is negative: $k\in i\mathbb{R}$ and the field solutions oscillate spatially inside the sphere. The factors $A^{in/out}_{S},\;B^{in/out}_{S}$ are merely constants of integration, which may be determined by requiring that the field be continuous and differentiable at $r\in\{0,\penalty 10000\ R\}$, and by imposing the boundary condition in Eq. (15). The solution is then:
$\varphi_{S}(r)=\begin{cases}\operatorname{sech}(kR)\frac{\sinh(kr)}{kr}&r<R\\ 1+\dfrac{q(kR)R}{r}&r>R\;,\end{cases}$ (18)
@@
$q(kR)=\frac{\tanh(kR)-kR}{kR}\penalty 10000\ .$ (19)
@@
$q(kR)\approx\begin{cases}-\frac{1}{3}(kR)^{2}&|kR|\ll 1\\ -1&|kR|\gg 1\;.\end{cases}$ (20)
The case $|kR|\ll 1$ is interpreted as a regime of weak coupling between sphere and field, in which the characteristic length scale of the field in matter, $1/|k|$, is considerably larger than the sphere. Hence, in this region, the field is minimally perturbed, and $\varphi_{S}\approx 1$. Conversely, in the case $|kR|\gg 1$, the sphere and field can be considered strongly coupled. As can be seen in Fig. (1), this leads to a general trend in which the field is increasingly screened in the vicinity of the sphere as $|kR|$ increases.
However, in the case $\alpha<0$, we observe additional phenomenology. Since $k\in i\mathbb{R}$, the $\tanh(kR)$ term in Eq. (19) now becomes $\sim\tan(|k|R)$. The field therefore diverges asymptotically at $|k|R=\pi\left(n+\frac{1}{2}\right)$, for integer $n$, and hence has an enhanced amplitude in the neighbourhood of these points. Whilst the condition $|\alpha\phi^{2}|\ll 1$ does not hold sufficiently close to these divergences, there may still be $|k|$ for which some degree of enhancement can be physical, depending on the value of $m$ associated with the field. Conversely, we note that there are additionally values for which $\tan(|k|R)=|k|R$, and so the effective scalar charge of the source vanishes, $q(kR)=0$, and the field profile around the sphere will be indistinguishable from the uncoupled case.
Figure 1: Evaluation of the field profile at the surface of a solid sphere radius $R$, varying $|\alpha\rho_{S}R^{2}|=|kR|^{2}$. The top panel illustrates the case $\alpha>0$, in which the field at the surface is screened for $\alpha=1/\rho R^{2}$. The lower panel demonstrates the case $\alpha<0$. Maxima and minima indicate the locations of field divergences and zeros respectively. We note that the magnitude of the field is displayed given that, in the $\alpha<0$ case, sign changes occur between adjacent extrema.
Additionally, for the case $\alpha<0$, it is necessary to acknowledge the restricted parameter space in which the solutions in Eq.(18) are stable. These solutions assume the field to have an expectation value $\left<\phi\right>=0$, both inside and outside matter. This is clearly a reasonable assumption if the mass-squared of the scalar is positive within the sphere: $m_{eff}^{2}=m^{2}+\rho\alpha>0$. Less trivially, it is also a valid in the case where the field is tachyonic inside the sphere: $m_{eff}^{2}=m^{2}+\rho\alpha<0$, but where $|k|R\ll 1$. Under these circumstances, the field expectation value has insufficient space inside the sphere to roll to a new minimum. In the case that both of these conditions are violated: $m^{2}\lesssim\rho|\alpha|$ and $|k|R\gtrsim 1$, we expect the field solutions in Eq. (18) to be unstable to perturbations. The value at which the field stabilises will be determined by terms beyond the quadratic cutoff of the field theory in Eq. (1). As such, we are unable to describe this region of parameter space - shown in Fig. 2 - in our scalar theory.
Figure 2: The region of instability of the scalar field solutions around a uniform sphere. The mass $m$ and coupling strength $\alpha$ of the scalar field are appropriately scaled according to the sphere’s density $\rho_{S}$ and radius $R$. The red dotted line illustrates the boundary below which the field is tachyonic. The black dotted line illustrates the boundary under which the wavelength of the field inside matter is sufficiently small to allow its expectation value to roll away from zero. The red shaded region illustrates the intersection of these regions: $\alpha<-\max{(\frac{m^{2}}{\rho},\frac{1}{\rho R^{2}})}$, whereby the field solutions in Eq. (18) are expected to be unstable to perturbations.
@@
$\varphi_{\oplus}(r)=\varphi_{S}(r\;;R_{\oplus},\penalty 10000\ k_{\oplus})\;,$ (21)
for which $k_{\oplus}=\sqrt{\rho_{\oplus}\alpha}$, $\rho_{\oplus}\simeq 5510\penalty 10000\ \mbox{kg m}^{-3}$ and $R_{\oplus}=6.4\times 10^{6}\penalty 10000\ m$.
@@ III.2 Spherical Cavity
where, $r,\penalty 10000\ \theta$ form part of the spherical co-ordinate basis about the cavity centre, defining $\theta\in\{0\leq\theta<\pi\}$ such that $\theta=0$ aligns with the direction of the field gradient. A caveat here is that, under our own description, the background field to the cavity should tend to the homogeneous solution away from the matter sourcing it. This requirement is in conflict with Eq. (24), which diverges as $r\rightarrow\infty$. However, we may discard this problem if the background field is sufficiently slowly varying, such that it can be described only by the leading order terms in its Taylor expansion in the vicinity of the cavity. These terms can be used to define $a_{0},\penalty 10000\ a_{1}$. For example, by expanding $\varphi_{\oplus}(r^{\prime};k_{\oplus},R_{\oplus})$, about the position of the cavity relative to the centre of the Earth, $\vec{r}_{c}$, we might define $a_{0},\penalty 10000\ a_{1}$ using the field sourced by the Earth:
@@
The validity of this interpretation when evaluating $a_{0},\penalty 10000\ a_{1}$ for different parameters is discussed in Sec. IV. With the inclusion of these boundary conditions, the system is no longer spherically symmetric and the equation of motion for the field is:
$\frac{1}{r^{2}}\partial_{r}(r^{2}\partial_{r}\varphi_{c})+\frac{1}{r^{2}\sin\theta}\partial_{\theta}\left(\sin\theta\partial_{\theta}\varphi_{c}\right)=\rho\alpha\varphi_{c}\;.$ (26)
By imposing the boundary conditions in Eq. (24), demanding continuity and differentiability of the field at the boundaries $r\in\{0,\penalty 10000\ R_{1},\penalty 10000\ R_{2}\}$, and through orthogonality of angular modes, solutions simplify to the form:
$\varphi_{c}(r)=\begin{cases}A_{0}^{in}+A_{1}^{in}r\cos\theta&r<R_{1}\\ \\ \begin{split}&\left(A_{0}^{walls}[j_{0}(-ikr)]+B_{0}^{walls}[(i)y_{0}(-ikr)]\right)\\ &+\left(A_{1}^{walls}[(-i)j_{1}(-ikr)]+B_{1}^{walls}[(-1)y_{1}(-ikr)]\right)\cos\theta\end{split}&R_{1}<r<R_{2}\\ \\ \left(a_{0}+B_{0}^{out}r^{-1}\right)+\left(a_{1}r+B_{1}^{out}r^{-2}\right)\cos\theta&r>R_{2}\;,\end{cases}$ (27)
where $A^{in/out}_{l},\penalty 10000\ B^{in/out}_{l}$ are constants of integration, $j_{l}(x),\penalty 10000\ y_{l}(x)$ are spherical Bessel functions of the first and second kind respectively, $k=\sqrt{\alpha\rho_{c}}$ and $i$ is the imaginary unit. Further detail on the origin of these terms, and analytical expressions for these coefficients are are presented in Appendix A.
We note that, as with the field solutions around a uniform sphere, there should be a region of parameter space similar to that in Fig. 2, under which the solutions defined by Eq. (27) are unstable to perturbations. In this work, the scale and density of the cavities considered will be smaller than the corresponding properties of Earth - which is modelled as a uniform sphere in Sec. III.1. As such, the region of instability for the field solutions around these cavities will be contained within the region of instability for field solutions around the Earth. For simplicity, we always choose to exclude the full region in which the field solutions around the Earth are unstable in our analyses.
@@
$a_{0}\rightarrow A^{in}_{0}$ $=a_{0}\times\frac{1}{kR_{1}\sinh(k(R_{2}-R_{1}))+\cosh(k(R_{2}-R_{1}))}\;,$ (28)
$a_{1}\rightarrow A^{in}_{1}$ $=a_{1}\times\frac{3kR_{2}}{3\sinh(k(R_{2}-R_{1}))+3kR_{1}\cosh(k(R_{2}-R_{1}))+(kR_{1})^{2}\sinh(k(R_{2}-R_{1}))}\;.$ (29)
@@
Figure 3: The modification of coefficients describing the field inside a spherical cavity: $R_{1}=10\penalty 10000\ \mbox{cm}$, $R_{2}=11\penalty 10000\ \mbox{cm}$ and $\rho_{c}=2700\penalty 10000\ \mbox{kg m}^{-3}$. Panels in the top row relate to coefficients for the positive $\alpha$ case, and the bottom row the negative $\alpha$ case. The left panels give the magnitude of the ratio between the constant component of the field profile inside the cavity, as in (27), and the boundary condition in Eq. (24). The right gives the same but for the linear component of the profile and boundary conditions. Note that the magnitude is plotted since for the $\alpha<0$ case, components of the field profile inside the cavity may change sign with respect to the boundary conditions. The blue dashed line indicates where $\alpha=k_{c}^{2}/\rho_{c}$.
@@
We demonstrate the effect of this modification in Fig. 3 for a small test cavity, inner radius $R_{1}=10\penalty 10000\ \mbox{cm}$ and outer radius $R_{2}=11\penalty 10000\ \mbox{cm}$, and with density similar to aluminium $\rho_{c}=2700\penalty 10000\ \mbox{kg m}^{-3}$. These dimensions have been chosen as to approximate the scale and properties of laboratory scale apparatus such as optical cavities. Other choices of size and material will exhibit qualitatively similar behaviour.
@@
As with the uniform spherical case, we may choose to broadly divide the solution for the field into two regimes, by comparing the value of $|k|$ to the thickness of the wall of the cavity. The critical value for a cavity is: $k_{c}=1/\left(R_{2}-R_{1}\right)$. For $|\alpha|\ll k_{c}^{2}/\rho_{c}$, such that the wavelength of the field in the cavity walls is larger than the cavity thickness $R_{2}-R_{1}$, Fig. 3 shows that the cavity has minimal impact on the amplitude of the field in the interior. Indeed, Eqs. (28), (29) demonstrate that $A_{0},\penalty 10000\ A_{1}$ should be well approximated by $a_{0},\penalty 10000\ a_{1}$ for $|k|\ll k_{c}$, aside for the region $1/R_{1}\lesssim|k|<k_{c}$, in which polynomial decay arising from the $kR_{1}$ terms in the denominators could be observed.